OpenStudy (anonymous):

A mixture of 30 lbs. of candy sells for $1.10 a pound. The mixture consists of chocolates worth $1.50 a pound and chocolates worth 90¢ a pound. How many pounds of each kind were used to make the mixture? ___ lbs. of $1.50 chocolates and ____ lbs. of 90¢ chocolate

OpenStudy (anonymous):

@Chlorophyll

OpenStudy (anonymous):

The concept is still MIXTURE, this time the mass multiply by its price ( instead of percentage) :)

OpenStudy (anonymous):

30x1.10+1.50x=(.90)x(30+x)

OpenStudy (anonymous):

15x + 9 ( 30 -x ) = 11 * 30

OpenStudy (anonymous):

confused

OpenStudy (anonymous):

Think about which one is product, then arrange it to the right side!

OpenStudy (anonymous):

" A mixture of 30 lbs. of candy sells for $1.10 a pound " is the product.

OpenStudy (anonymous):

ok 30x11right

OpenStudy (anonymous):

110 cents * 30 lbs However in my equation, I already simplified all by 10: -> 11 * 30